Weizmann Workshop on Multi Party Secure Computation

Weizmann Institute of Science, Rehovot, Israel

June 16-18, 1998

Workshop Committee  Workshop Coordinator
Shafi Goldwasser  Carol Weintraub
Moni Naor  Phone: (972) (8) 934-3545

Fax: (972) (8) 934-4122

Email: carol@wisdom.weizmann.ac.il

Location

The Workshop will be held Ziskind Building, Room number 1, at the Weizmann Institute of Science, Rehovot, Israel. Phone number at workshop location is (972) (8) 934-3540

The Weizmann campus is in Rehovot, a town located about 15 miles (25km) from Tel-Aviv, and 35 miles (55km) from Jerusalem. Buses are available to both cities, but private transportation is much easier.

You are most likely to arrive in Israel via the Ben-Gurion International Airport (sometimes referred to as the “Tel-Aviv Airport”^{1}). This airport is about 10 miles from Rehovot, 15 miles from Tel-Aviv, and 30 miles from Jerusalem.

Registration

There is no registration fee. However, we do ask that you complete the enclosed questionnaire and return it to us as soon as possible (even with partial information to be complete at a later time). A registration package will be waiting for you in the lobby of Ebner Hall, containing a name tag. Name tags will allow access to meals during the workshop.

1which is a bit confusing, since Tel-aviv also has a small airport within the city limits.

Accommodations

A block of rooms has been reserved on the Weizmann campus for the workshop period. Most rooms are available starting Tuesday, June 15th, till Friday, June 19th. A few are also available for Friday night. We offer to help you with hotel reservations in Tel-Aviv or Jerusalem for the time beyond the workshop, and in particular for Wednesday night. See details below.

The rooms are at: the youth center (known as “Mer-kaz Ha-no-ar”), and a couple (one or two) at San-Martin. The youth centers has no phones in the rooms, but there are public phones outside. These rooms seem to be in high demand throughout the institute, and so we were not able to obtain an optimal arrangement (such as allowing each participant to have a room by him/herself for any period around the workshop he/she prefers). There is some possibility that a few participants will have to share rooms in order to accommodate everyone on campus. Thus, we are asking you if you are willing to share a double room (and with whom).

After the Conference

To obtain a reasonable international air-fare ordinarily one must stay in Israel at least six nights. Furthermore, you will probably have to stay in Israel on Thursday night if you intend to attend the technical sessions on that day.

The conference coordinator, Ms. Carol Weintraub, will be happy to help you with hotel reservations in either cities, if you let her know in advance. In case you are looking for an inexpensive hotel (say, around $50 per night), may we recommend the Atlas Business Club in Tel-Aviv which owns several inexpensive hotels around Tel-Aviv (972-3-546-8126), Hotel Carlton in Tel-Aviv (972-3-520-0278), or Hotel Renaissance in Jerusalem (972-2-528-111).

Arriving

On the day and night of June 15th and/or June 16th (depending when you plan to arrive). there will be someone at the gate of Weizmann institute who will have the key to your room and an envelope with all necessary information to direct you to your room and the workshop. If you are arriving on a different date, please make arrangements with Ms. Carol Weintraub. Arriving by taxi: Upon arriving to Ben-Gurion Airport, and once out of customs, directly outside the gate is an airport taxi station. Ask for a taxi to the Weizmann Institute, Rehovot. The fare should run around $15 (45 NIS). At the main entrance gate to the Weizmann Institute have the taxi driver ask the guard for an envelope intended for you (if you stay on campus), and how to arrive to your room which will be at the youth center (known as “Mer-kaz Ha-no-ar”) most likely, or the San Martin Guest House. Arriving by rental car: You can arrange a rental car to be picked up at the airport (or the big cities) from all the usual companies (Avis, Hertz, Budget). However, the best fare is usually offered by an Israeli company called Eldan which has an office in Rehovot as well. Furthermore, Eldan offers a 10% reduction to the visitors of Weizmann Institute and it has been confirmed that you are entitled to this reduction (please notify Carol in case of trouble!). The numbers of Eldan are 972-8-468-376 (in Rehovot), 1-212-629-6090 (in NYC), and 1-800-938-5000 (toll free in the States). You may get a better rate if you reserve the car from the States. If you rent a car, you should ask the rental company for a map and exact instructions how to get to the Weizmann Institute in Rehovot. Getting to Rehovot from the Tel-Aviv – Jerusalem Road (on which also the Ben-Gurion Airport lies) is not hard (you just have to take three correct turns), and once you arrive in Rehovot you cannot possibly miss the Weizmann Institute (which is located on the main road which enters the town, just 200 meters after a railroad crossing). Parking: Parking will be available both for those staying on campus, and for lecture attendees coming from outside the Institute.

Climate

The month of June is usually hot during the day and hot during the night. Yet, it is even hotter in July and August, in which we would have said ”unbearably hot”. In Jerusalem the nights are mild, you should bring a sweater, although I doubt you would use it.

Jerusalem  Tel-Aviv and Rehovot
18^{∘}–27.5^{∘}C 20^{∘}–27.2^{∘}C
64.4^{∘}–81.5^{∘}F 68^{∘}–80.9^{∘}F

Meyer & Shirley Weisgal Recreation Center

The Weisgal Recreation Center, situated on campus—with indoors swimming pool, tennis courts, sauna and fitness gymnasium—is open from 6:00–18:00. Participants and accompanying persons may use the facilities upon presentation of their name badge.

Visa

Some nationalities require a visa to enter Israel (e.g. India, Russia). Please check if you require a visa. If a visa is required, let us know and we will mail or fax you a formal letter of invitation to present to the Israeli consulate.

Important Numbers and Addresses

• Prof. Shafi Goldwasser, Workshop Organizer, Dept. of Applied Mathematics and Computer Science, Weizmann Institute of Science, Rehovot, Israel. Tel: (972) (8) 9342953, Fax: (972) (8) 9344122,

• Prof. Moni Naor, Workshop Organizer, Dept. of Applied Mathematics and Computer Science, Weizmann Institute of Science, Rehovot, Israel. Tel: (972) (8) 9343701, Fax: (972) (8) 9344122,

Email: shafi@wisdom.weizmann.ac.il.

• Ms. Carol Weintraub, Workshop Coordinator, Dept. of Applied Mathematics and Computer Science, Weizmann Institute of Science, Rehovot, Israel. Tel: (972) (8) 9343545, Fax: (972) (8) 9344122, Email: carol@wisdom.weizmann.ac.il.

• Mr. Raanan Michael, Chief Administrator of the Faculty of Mathematical Sciences, Weizmann Institute of Science, Rehovot, Israel. Tel: (972) (8) 9343758, Fax: (972) (8) 9344122, Email: raanan@wisdom.weizmann.ac.il.

• The Hermann Mayer French House (San Martin)

Telephone: (972) (8) 9343385, Fax: (972) (8) 9342431

==========================================================================
========      W O R K S H O P       Q U E S T I O N N A I R E     ========
==========================================================================
Please fill-up and return this questionnaire ASAP
to Mrs Carol Weintraub (the workshop coordinator)
at email carol@wisdom.weizmann.ac.il
or fax to (972) (8) 344122.
FAMILY NAME:
FIRST NAME:
AFFILIATION:
EMAIL:
PHONE:
ANY SPECIAL REQUESTS:

ARRIVAL/DEPARTURE INFORMATION:
------------------------------
Accurate information is important in case you wish to have accommodations
on campus, but also otherwise it will be nice to know when to expect you.
IN CASE YOUR PLANS ARE NOT FIXED YET, PLEASE COMPLETE WHATEVER YOU KNOW
AND UPDATE CAROL (SEE ABOVE) ON THE MISSING DETAILS AT A LATER TIME.
DATE OF ARRIVAL:
EXPECTED TIME OF ARRIVAL:
DATE OF DEPARTURE:
EXPECTED TIME OF DEPARTURE:
DO YOU NEED A ROOM ON CAMPUS: yes/no
WOULD YOU LIKE THE ROOM ALSO FOR WEDNESDAY NIGHT:
ANY ACCOMPANYING PERSONS (IF SO SUPPLY DETAILS):
ARE YOU WILLING TO SHARE A ROOM: yes(with whom)/no
Do you want our help with hotel reservations in Tel-Aviv or Jerusalem
for the period beyond the workshop and in particular for Wednesday night)?
If so please see the "beyond the workshop" section in the following
Additional Information sheet and provide us with your preferences.